# Equilibrium-distribution-function based mesoscopic finite-difference methods for partial differential equations: Modeling and Analysis

### v1

@@ Abstract
In this paper, based on the idea of direct modeling with equilibrium distribution functions (EDFs) we develop a general framework of mesoscopic numerical method (MesoNM) for macroscopic partial differential equations (PDEs), including but not limited to nonlinear convection-diffusion equation (NCDE) and the Navier-Stokes equations (NSEs). Different from the mesoscopic lattice Boltzmann method, this kind of MesoNM is an EDF-based mesoscopic finite-difference (MesoFD) method, and through taking the moments of MesoFD scheme its macroscopic version, called MMFD method, can be derived directly. Both MesoFD scheme and MMFD schemes are muti-level FD methods, MesoFD scheme being mesoscopic, and MMFD scheme being its macroscopic form which has the form of the central FD scheme. They are unified FD schemes for PDEs and can be implicit or explicit forms as needed. The macroscopic moment equations (MEs) can be derived from the MesoFD or MMFD scheme through the Taylor expansion method, and the common PDEs can be recovered from the MEs by using direct Taylor expansion method. Moreover, the stability of MMFD scheme is analyzed for linear CDE with anisotropic diffusion, and the stability conditions of a two-level explicit MMFD scheme and a two-level $\theta$-MMFD scheme (hybrid explicit and implicit MMFD scheme) are obtained, respectively. Finally, we note that some existing lattice Boltzmann (LB) based macroscopic FD models for the NSEs and NCDE are the special cases of present MMFD, which can be considered as a unified framework of FD schemes for PDEs, from this point of view.
@@ I Introduction
The lattice Boltzmann (LB) method, as an effective mesoscopic numerical method (MesoNM) based on kinetic theory, has gained great success in the study of complex flows, heat and mass transfer described by the Navier-Stokes equations (NSEs), nonlinear convection-diffusion equations (NCDEs), and the coupled systems of them [1, 2, 3, 4]. Up to now, although there are many different types of LB models [16], the most popular ones are the single-relaxation-time LB (SRT-LB) model or lattice Bhatnagar-Gross-Krook (LBGK) model [5, 6], the two-relaxation-time LB (TRT-LB) model [7, 8, 9, 10, 11], and the general or multple-relaxation-time LB (MRT-LB) model [12, 13, 14, 15]. The MRT-LB model has multiple relaxation parameters, and compared to the SRT version, it can achieve more stable and more accurate results through adjusting the free relaxation parameters [12, 13, 14, 15], also the TRT-LB model with only one free relaxation parameter [7, 8, 9, 10, 11]. Actually, the classic MRT-LB model can be more general through introducing a block-lower-triangular-relaxation matrix [16], and rectangular grid [17].
@@
It is known that the LB method is based on the evolution of the distribution function (DF) in all lattice velocity directions at each lattice node, and the macroscopic conserved quantities are computed from the moments of the DF. This evolution based on the DF brings us several challenging problems that need to be addressed: (1) a large amount of memory is required; (2) the initial and boundary conditions for macroscopic quantities should be treated carefully; (3) more importantly, the theoretical analysis (e.g., stability analysis) on the LB method is complicated and difficult. To overcome these problems, it is desirable to explore the connection between LB method and the traditional numerical method. In the past years, some efforts have been made to derive the macroscopic numerical scheme from the LB method.
@@
In the LB community, it is well known that the SRT-LB model with relaxation time $\tau=1$ can be written as a macroscopic FD scheme [18, 19, 20, 21]. In order to overcome the limitation $\tau=1$, the lattice kinetic scheme was proposed in Ref. [20], and several modified versions have been developed in the literature. In addition to those derived from the SRT-LB model [18, 19, 20, 21], up to now there have been many research works on macroscopic FD schemes derived or developed from LB models, e.g., the TRT-LB model [9, 22, 10], and some other type of macroscopic FD schemes, such as the so-called simplified lattice Boltzmann method [23, 24] and its modified models [25, 26], and the recursive FD LB models [27, 28]. In addition, for the one-dimensional case, some three- and four-level FD schemes can also be obtained from the LB models [29, 30, 31]. It should be noted that based on the recurrence equations with $s^{+}+s^{-}=2$, a class of three-level FD schemes can be derived from TRT-LB model [9, 22, 10], which contains the FD scheme obtained by the SRT-LB model as its special case (setting $s^{+}=s^{-}=1$), where $s^{+}$ and $s^{-}$ are the relaxation parameters corresponding to the even-order and odd-order moments, respectively.
@@
It can be seen that the above works on the macroscopic FD schemes [32] derived from the LB models are usually limited to special cases. Recently, Bellotti et al. [33, 34] demonstrated that any LB scheme can be rewritten as a corresponding multi-level FD scheme on the conserved variables, and conducted a rigorous consistency study and the derivation of the modified equations. However, as stated in Ref. [33], the multi-level FD scheme usually has the theoretical value due to its complicated multi-level structure and the formalism based on operators, commutative algebra and polynomials (’we stress once more the fact that recasting a lattice Boltzmann scheme as a Finite Difference scheme on the conserved moments is merely useful from the theoretical perspective, being of moderate interest for actual efficient implementations’ [33]). Based on the results in Refs. [33, 34], some specific LB models and corresponding multi-level FD schemes with high-order accuracy for one- or two-dimensional simple problems have been designed [31, 35]. According to the previous work [16], Chen et al. [36] presented a general propagation multiple-relaxation-time lattice Boltzmann (GPMRT-LB) model, and then following the works [33, 34], they obtained the macroscopic FDS and modified equations corresponding to the GPMRT-LB model.
@@
Different from the above methods for deriving the FD schems from the existing LB models, in this work, we uses the equilibrium distribution function (EDF) to represent the distribution function (DF), and propose a novel class of mesoscopic numerical methods (MesoNMs), called the mesoscopic finite-difference (MesoFD) schemes, for partial differential equations (PDEs) with the source terms. Then the MesoFD based macroscopic finite difference (MMFD) schemes can be directly obtained by taking the moments of the MesoFD scheme. Both of them are of multi-level implicit or explicit structure. The MMFD scheme can be viewed as a unified framework of FD schemes for PDEs, and the commonly used FD schemes derived from LB models and traditional FD schemes are its special cases.
@@ II Mesoscopic finite difference method and macroscopic finite difference scheme
From Eq. (1) we can also obtain the evolution equations of required moments. Similar to the LB method, the unknown macroscopic conserved variable(s), that is, the zeroth moment $\phi(\mathbf{x},t)$ for scalar PDE, such as NCDE, or the zeroth moment $\rho(\mathbf{x},t)$ and the first moment $(\rho\mathbf{u})(\mathbf{x},t)$ for NSEs, can be computed by
@@
in which two different sets of parameters, $\{a_{kq},b_{kq}\}$ and $\{\bar{a}_{kq},\bar{b}_{kq}\}$ are used.
Remark 1 We can replace $f_{j}$ and $a_{00}$ in Eq. (1) by $f_{j}^{ne}$ and $a_{00}-1$, respectively, then we obtain the equivalent form of Eq. (1). Therefore, if the summation is only made for the case $k=q$, one can easily find that the scheme in Ref. [28] is just a special case of the MesoFD scheme (1). Unless otherwise stated, the following discussion is about Eq. (1).
@@ II.1 The central difference representation of MesoFD method
Interestingly, both MesoFD and MMFD schemes can be represented by the central difference format. First, we define the following central difference operators through introducing a function $g(\mathbf{x})$, and two integers $j$ and $k$,
@@
Using the above operators, one can easily obtain the central difference representation of MesoFD scheme (1):
@@
where $S,\mathbf{F}$ in Eqs. (16) and (17) and the parameters in Eq. (17b) are defined as $S=\sum_{j}F_{j}=\sum_{j}F_{j}^{+},~{}\mathbf{F}=\sum_{j}\mathbf{c}_{j}F_{j}=\sum_{j}\mathbf{c}_{j}F_{j}^{-},~{}$ $\bar{a}_{kq}^{+}=\bar{a}_{kq}+\bar{a}_{-kq},~{}\bar{a}_{kq}^{-}=\bar{a}_{kq}-\bar{a}_{-kq},~{}\bar{b}_{kq}^{+}=\bar{b}_{kq}+\bar{b}_{-kq},~{}\bar{b}_{kq}^{-}=\bar{b}_{kq}-\bar{b}_{-kq}$, and $\bar{a}_{kq},\bar{b}_{kq}$ are used as those in Eq. (6). If needed, one can take the higher-order moments in Eq. (13) to get the evolution equations of these higher-order moments.
It can be seen from Eqs. (16) and (17) that the MMFD scheme consists of central difference schemes on multiple-time levels. Based on the MesoFD scheme (Eq. (1)) and the MMFD scheme (Eq. (16) or Eq. (17)), we can derive the macroscopic PDEs from mesoscopic and macroscopic ways, respectively.
@@ II.2 The two-level central MMFD scheme
As we know, the two-level FD scheme is the most commonly used one for its efficiency . In this part, we only consider the two-level central MMFD scheme for NCDE, and the two-level central MMFD scheme for NSEs can be given similarly. Taking $q=-1$ in Eq. (16), we obtain the explicit MMFD (eMMFD) scheme
@@
and taking $q\in\{0,-1\}$ in Eq. (16), the implicit MMFD (iMMFD) scheme can be obtained as follows
@@ III The moment equations of MesoFD method: Direct Taylor expansion
There are four basic analysis methods that can be used to recover the macroscopic NSEs and NCDE from the LB models, i.e., the Chapman-Enskog (CE) analysis [37], the Maxwell iteration (MI) method [38], the direct Taylor expansion (DTE) method [16] and the recurrence equations (RE) method [10, 9, 11]. In Ref. [16], these four methods have been compared, and it is shown that they can give the same equations at the second-order of expansion parameters, while the DTE method is much simpler. The MI method and RE method may not be suitable for analyzing the MesoFD scheme. In what follows, the DTE method is used to analyze the MesoFD scheme at the acoustic scaling.
@@ III.1 The derivation of the moment equation for PDE using DTE method
Taking $A_{00}=1$ in Eq. (III.1), we obtain a general moment equation for PDE of order $N$,
@@
We note that Eq. (31c) is the moment equation corresponding to a third-order PDE, while Eq. (31b) is a second-order approximation and is also a moment equation for second-order PDE. Eq. (31a) can be taken as their first-order approximation. Some popular second-order PDEs, such as NCDE (DE), convection equation, Poisson equation, and even wave equation can be derived from Eq. (31b) or Eq. (31c) when $A_{lm},B_{lm},\mathbf{\mathbf{M}}_{k}$, and $\mathbf{M}_{kF}$ are properly given.
@@
then we obtain $-\frac{B_{00}}{A_{10}}M_{0F}=S$, and Eq. (33) becomes
@@
Set $M_{0}=0$ or consider a steady case, Eq. (31b) becomes
@@
Taking $\mathbf{M}_{1}=\mathbf{B},~{}\mathbf{M}_{2}=\beta c_{s}^{2}\mathbf{D},~{}M_{0F}=S,~{}\mathbf{M}_{1F}=\mathbf{0},~{}B_{00}=-A_{11}$, and $\alpha=-\frac{A_{22}}{A_{11}}\Delta t\beta c_{s}^{2}$ in Eq. (42), the steady CDE is recovered
@@
Let $A_{20}=-1,~{}A_{lm}=0$ for $(l,m)\neq(2,0)$ and $(2,2)$, and $B_{lm}=0$, for all $l,m$ in Eq. (31c) a wave equation is recovered as follows.
@@
Taking $M_{0}=\phi,~{}\mathbf{M}_{2}=\beta c_{s}^{2}\mathbf{D}$ in Eq. (44), we obtain the following wave equation
@@
Remark 1: (1) When $\mathbf{M}_{1}=\mathbf{B}=\mathbf{0}$, Eq. (41) becomes a diffusion equation, and we choose $\mathbf{C}=0$, thus $\mathbf{M}_{1F}=0$. In this case, there is no restriction on $A_{11}$, and only $A_{10}\neq 0$ is needed.
(2) The source term $S=0$ or $S$ is a nonzero constant, then Eq. (38) is satisfied. Otherwise, $B_{10}+A_{20}=0$ is needed.
(3) When $\mathbf{D}=\mathbf{0}$, Eq. (41) becomes a convection equation (CE), which means that the CE can also be recovered directly from the MesoFD method. Another way to derive CE from the MesoFD method is to modify the diffusion coefficient $\alpha$ as a small quantity of order $\Delta t$ or $\Delta x$.
(4) If $\mathbf{M}_{1}=0$, and $\mathbf{M}_{3}=0$ in Eq. (31c), then to obtain Eq. (44), $A_{11},A_{21}$ and $A_{33}$ can be non-zero.

### v2

@@ Abstract
In this paper, based on the idea of direct discrete modeling (DDM) with equilibrium distribution functions (EDFs), we develop a general framework of the mesoscopic numerical method (MesoNM) for macroscopic partial differential equations (PDEs), including but not limited to the nonlinear convection-diffusion equation (NCDE) and the Navier-Stokes equations (NSEs). Unlike the mesoscopic lattice Boltzmann method, this kind of MesoNM is an EDF-based mesoscopic finite-difference (MesoFD) method, and by taking the moments of the MesoFD scheme, its macroscopic version, called MMFD method, can be derived directly. Both MesoFD scheme and MMFD schemes are multi-level FD methods, MesoFD scheme being mesoscopic, and MMFD scheme being its macroscopic form which has the form of the central FD scheme. They are unified FD schemes for PDEs and can be in implicit or explicit forms as needed. The macroscopic moment equations (MEs) can be derived from the MesoFD or MMFD scheme through the Taylor expansion method, and the common PDEs can be recovered from the MEs by using the direct Taylor expansion method. Moreover, the stability of the MMFD scheme is analyzed for linear CDE and liner wave equation with anisotropic diffusion, and the stability conditions of a two-level explicit MMFD scheme, a two-level $\theta$-MMFD scheme (hybrid explicit and implicit MMFD scheme), and a three-level MMFD scheme are obtained, respectively. Finally, we note that some existing lattice Boltzmann (LB) based macroscopic FD models for the NSEs and NCDE are the special cases of present MMFD, which can be considered as a unified framework of FD schemes for PDEs, from this point of view.
@@ I Introduction
The lattice Boltzmann (LB) method, as an effective mesoscopic numerical method (MesoNM) based on kinetic theory, has gained great success in the study of complex flows, heat and mass transfer described by the Navier-Stokes equations (NSEs), nonlinear convection-diffusion equations (NCDEs), and their coupled systems [1, 2, 3, 4]. Until now, although there are many different types of LB models [16], the most popular are the single-relaxation-time LB (SRT-LB) model or lattice Bhatnagar-Gross-Krook (LBGK) model [5, 6], the two-relaxation-time LB (TRT-LB) model [7, 8, 9, 10, 11], and the general or multple-relaxation-time LB (MRT-LB) model [12, 13, 14, 15]. The MRT-LB model has multiple relaxation parameters, and compared to the SRT version, it can achieve more stable and more accurate results by adjusting the free relaxation parameters [12, 13, 14, 15], also the TRT-LB model with only one free relaxation parameter [7, 8, 9, 10, 11]. In fact, the classic MRT-LB model can be more general by introducing a block-lower-triangular-relaxation matrix [16], and rectangular grid [17].
@@
It is known that the LB method is based on the evolution of the distribution function (DF) in all lattice velocity directions at each lattice node, and the macroscopic conserved quantities are computed from the moments of DF. This evolution based on the DF brings us several challenging problems that need to be addressed: (1) a large amount of memory is required; (2) the initial and boundary conditions for macroscopic quantities should be treated carefully; (3) more importantly, the theoretical analysis (e.g., stability analysis) on the LB method is complicated and difficult. To overcome these problems, it is desirable to explore the connection between the LB method and the traditional numerical method. In the past years, some efforts have been made to derive the macroscopic numerical scheme from the LB method.
@@
In the LB community, it is well known that the SRT-LB model with relaxation time $\tau=1$ can be written as a macroscopic FD scheme [18, 19, 20, 21]. In order to overcome the limitation $\tau=1$, the lattice kinetic scheme was proposed in Ref. [20], and several modified versions have been developed in the literature. In addition to those derived from the SRT-LB model [18, 19, 20, 21], up to now there have been many research works on macroscopic FD schemes derived or developed from LB models, e.g., the TRT-LB model [9, 22, 10], and some other type of macroscopic FD schemes, such as the so-called simplified lattice Boltzmann method [23, 24] and its modified models [25, 26], and the recursive FD LB models [27, 28]. Furthermore, for the one-dimensional case, some three- and four-level FD schemes can also be obtained from the LB models [29, 30, 31]. However, it should be noted that based on the recurrence equations with $s^{+}+s^{-}=2$, a class of three-level FD schemes can be derived from TRT-LB model [9, 22, 10], which contains the FD scheme obtained from the SRT-LB model as its special case (setting $s^{+}=s^{-}=1$), where $s^{+}$ and $s^{-}$ are the relaxation parameters corresponding to the even-order and odd-order moments, respectively.
@@
It can be seen that the above works on the macroscopic FD schemes [32] derived from the LB models are usually limited to some special cases. Recently, Bellotti et al. [33, 34] demonstrated that any LB scheme can be rewritten as a corresponding multi-level FD scheme on the conserved variables, and conducted a rigorous consistency study and the derivation of the modified equations. However, as stated in Ref. [33], the multi-level FD scheme usually has the theoretical value due to its complicated multi-level structure and the formalism based on operators, commutative algebra and polynomials (’we stress once more the fact that recasting a lattice Boltzmann scheme as a Finite Difference scheme on the conserved moments is merely useful from the theoretical perspective, being of moderate interest for actual efficient implementations’ [33]). Based on the results in Refs. [33, 34], some specific LB models and corresponding multi-level FD schemes with high-order accuracy for one- or two-dimensional simple problems have been designed [31, 35]. According to the previous work [16], Chen et al. [36] presented a general propagation multiple-relaxation-time lattice Boltzmann (GPMRT-LB) model, and then following the works [33, 34], they obtained the macroscopic FD schemes and modified equations corresponding to the GPMRT-LB model.
@@
Differently from the above methods for deriving the FD schemes from the existing LB models, in this work, we use the equilibrium distribution function (EDF) to represent the distribution function (DF), and propose a novel class of mesoscopic numerical methods (MesoNMs), called the mesoscopic finite-difference (MesoFD) schemes, for partial differential equations (PDEs) with the source terms. Then the MesoFD based macroscopic finite-difference (MMFD) schemes can be directly obtained by taking the moments of the MesoFD scheme. Both of them are of multi-level implicit or explicit structure. The MMFD scheme can be viewed as a unified framework of FD schemes for PDEs, and the commonly used FD schemes derived from LB models and traditional FD schemes are its special cases.
@@ II Mesoscopic finite difference method and macroscopic finite difference scheme
From Eq. (1) we can also obtain the evolution equations of required moments. Similar to the LB method, the unknown macroscopic conserved variable(s), for instance, the zeroth moment $\phi(\mathbf{x},t)$ for scalar PDE, such as NCDE, or the zeroth moment $\rho(\mathbf{x},t)$ and the first moment $(\rho\mathbf{u})(\mathbf{x},t)$ for NSEs, can be computed by
@@
in which two different sets of parameters, $\{a_{kq},b_{kq}\}$ and $\{\bar{a}_{kq},\bar{b}_{kq}\}$, are used.
Remark 1 We can replace $f_{j}$ and $a_{00}$ in Eq. (1) by $f_{j}^{ne}$ and $a_{00}-1$, respectively, then we obtain the equivalent form of Eq. (1). Therefore, if the summation is only made for the case $k=q$, one can easily find that the scheme in Ref. [28] is just a special case of the MesoFD scheme (1). Unless otherwise stated, the following discussion is only limited to Eq. (1).
@@ II.1 The central difference representation of MesoFD method
It is interesting that both MesoFD and MMFD schemes can be represented by the central difference format. First, we define the following central difference operators through introducing a function $g(\mathbf{x})$, and two integers $j$ and $k$,
@@
Using the above operators, one can obtain the central difference representation of MesoFD scheme (1):
@@
where $S,\mathbf{F}$ in Eqs. (16) and (17) and the parameters in Eq. (17b) are defined as $S=\sum_{j}F_{j}=\sum_{j}F_{j}^{+},~{}\mathbf{F}=\sum_{j}\mathbf{c}_{j}F_{j}=\sum_{j}\mathbf{c}_{j}F_{j}^{-},~{}$ $\bar{a}_{kq}^{+}=\bar{a}_{kq}+\bar{a}_{-kq},~{}\bar{a}_{kq}^{-}=\bar{a}_{kq}-\bar{a}_{-kq},~{}\bar{b}_{kq}^{+}=\bar{b}_{kq}+\bar{b}_{-kq},~{}\bar{b}_{kq}^{-}=\bar{b}_{kq}-\bar{b}_{-kq}$, $\bar{a}_{kq}and\bar{b}_{kq}$ are used as those in Eq. (6). If needed, one can take the higher-order moments in Eq. (13) to get the evolution equations of these higher-order moments.
It can be seen from Eqs. (16) and (17) that the MMFD scheme consists of central difference schemes at multiple time levels. Based on the MesoFD scheme (Eq. (1)) and the MMFD scheme (Eq. (16) or Eq. (17)), we can derive the macroscopic PDEs from mesoscopic and macroscopic ways, respectively.
It is worth noting that both Eq. (16) and Eq. (17) have the seperated form of spatio-temporal difference. Through using a divide-and-conquer strategy again, we can extend the one-set-parameter scheme Eq. (16) to a multiple-set-parameter one,
$\phi(\mathbf{x},t)$ $=$ $\sum_{k,q}\big{[}a_{kq}\phi+\Delta tb_{kq}S\big{]}(\mathbf{x},t+q\Delta t)+\frac{1}{2}\sum_{q}\sum_{j}\sum_{k=1}^{m_{q}}\big{[}\bar{a}_{kq}^{+}\Delta_{j,k}^{(2)}f_{j}^{eq,+}+\Delta t\bar{b}_{kq}^{+}\Delta_{j,k}^{(2)}F_{j}^{+}\big{]}(\mathbf{x},t+q\Delta t)$ (18)
$+\frac{1}{2}\sum_{q}\sum_{j}\sum_{k=1}^{m_{q}}\big{[}\hat{a}_{kq}^{-}\Delta_{j,k}^{(1)}f_{j}^{eq,-}+\Delta t\hat{b}_{kq}^{-}\Delta_{j,k}^{(1)}F_{j}^{-}\big{]}(\mathbf{x},t+q\Delta t),$
and correspondingly, the MesoFD scheme (1) has the following form,
$f_{j}(\mathbf{x},t)=\sum_{k,q}\big{[}\bar{a}_{kq}f_{j}^{eq[k,q],+}+\hat{a}_{kq}f_{j}^{eq[k,q],-}+\Delta t(\bar{b}_{kq}F_{j}^{[k,q],+}+\hat{b}_{kq}F_{j}^{[k,q],-})\big{]},$ (19)
where $\{a_{kq},b_{kq}\},\{\bar{a}_{kq},\bar{b}_{kq}\}$ and $\{\hat{a}_{kq},\hat{b}_{kq}\}$ can be different. Similarly, one can give the multiple-set-parameter version of scheme Eq. (17). It should be noted that Eq. (19) is a TRT-version of the MesoFD scheme (1), and one can also give a MRT-version of the MesoFD scheme (1) by assigning different coefficients to each term of $f_{j}^{eq}$. For simplicity, however, in the following, we will not discuss the TRT- and MRT- versions.
@@ II.2 The two-level central MMFD scheme
As we know, the two-level FD scheme is the most widely used one for its efficiency. In this part, we only consider the two-level central MMFD scheme for NCDE, and the two-level central MMFD scheme for NSEs can be given in a similar way. Taking $q=-1$ in Eq. (16), we obtain the explicit MMFD (eMMFD) scheme,
@@
and taking $q\in\{0,-1\}$ in Eq. (16), the implicit MMFD (iMMFD) scheme can be obtained as follows,
@@ III The moment equations of MesoFD method: Direct Taylor expansion
There are four basic analysis methods that can be used to recover the macroscopic NSEs and NCDE from the LB models, i.e., the Chapman-Enskog (C-E) analysis [37], the Maxwell iteration (MI) method [38], the direct Taylor expansion (DTE) method [16] and the recurrence equations (RE) method [10, 9, 11]. In Ref. [16], it is shown that they can give the same equations at the second-order of expansion parameters, while the DTE method is much simpler. The MI method and RE method may not be suitable for analyzing the MesoFD scheme. In what follows, the DTE method is used to analyze the MesoFD scheme at the acoustic scaling.
@@ III.1 The derivation of the moment equation for PDE using DTE method
Taking $A_{00}=1$ in Eq. (III.1), we can obtain a general moment equation for PDE of order $N$,
@@
We note that Eq. (33c) is the moment equation corresponding to a third-order PDE, while Eq. (33a) and Eq. (33b) are the first- and second-order approximation, and is also a moment equation for second-order PDE. Some popular second-order PDEs, such as NCDE, convection equation (CE), Poisson equation, and even wave equation (WE) can be derived from Eq. (33b) or Eq. (33c) when $A_{lm},B_{lm},\mathbf{\mathbf{M}}_{k}$, and $\mathbf{M}_{kF}$ are properly given.
@@
then we can obtain $-\frac{B_{00}}{A_{10}}M_{0F}=S$, and Eq. (35) becomes
@@
Let $M_{0}=0$ or consider a steady case, Eq. (33b) becomes
@@
Taking $\mathbf{M}_{1}=\mathbf{B},~{}\mathbf{M}_{2}=\beta c_{s}^{2}\mathbf{D},~{}M_{0F}=S,~{}\mathbf{M}_{1F}=\mathbf{0},~{}B_{00}=-A_{11}$, and $\alpha=-\frac{A_{22}}{A_{11}}\Delta t\beta c_{s}^{2}$ in Eq. (44), the steady CDE can be recovered
@@
Let $A_{20}=-1,~{}A_{lm}=0$ for $(l,m)\neq(2,0)$ and $(2,2)$, and $B_{lm}=0$, for all $l,m$ in Eq. (33c) a wave equation can be recovered
@@
Taking $M_{0}=\phi$ and $\mathbf{M}_{2}=\beta c_{s}^{2}\mathbf{D}$ in Eq. (46), we obtain the following wave equation,
@@
Remark 1: (1) When $\mathbf{M}_{1}=\mathbf{B}=\mathbf{0}$, Eq. (43) becomes a diffusion equation, and we can choose $\mathbf{C}=0$, thus $\mathbf{M}_{1F}=0$. In this case, there is no restriction on $A_{11}$, and only $A_{10}\neq 0$ is needed.
(2) If the source term $S=0$ or $S$ is a nonzero constant, then Eq. (40) is satisfied. Otherwise, $B_{10}+A_{20}=0$ is needed.
(3) When $\mathbf{D}=\mathbf{0}$, Eq. (43) becomes a CE, which means that the CE can also be recovered directly from the MesoFD method. Another way to derive CE from the MesoFD method is to modify the diffusion coefficient $\alpha$ as a small quantity of order $\Delta t$ or $\Delta x$, such as Lax-Wendroff scheme.
(4) If $\mathbf{M}_{1}=0$ and $\mathbf{M}_{3}=0$ in Eq. (33c), then to obtain Eq. (46), $A_{11},A_{21}$ and $A_{33}$ can be non-zero.